OpenStudy (anonymous):

lim as x->0+ of (1+2/x)^x

OpenStudy (anonymous):

i took the natural log on both sides and then rewrote to get infinity/infinity and then used L'Hopital's Rule, but I am having issues simplifying...

hartnn (hartnn):

so you got till here ? \(\large \ln L = \lim \limits_{x\to0^+ } x \ln (1+2/x)\) right?? then i guess you wrote that as \(\large \ln L = \lim \limits_{x\to0^+ } \dfrac{\ln (1+2/x)}{\dfrac{1}{x}}\) and applied LH rule ? what did you get after derivating the numerator and denominator ?

hartnn (hartnn):

for the derivative of numerator, you'll need CHAIN rule!

OpenStudy (anonymous):

that's where it got really messy. i got:  and (if that is even right) i have no idea where to go from there...)